Problem 5.1: Find the GS for y + y + y + y = 1 + e* + e-* + e2x + e-2x Problem 5.2: Find the GS for y(4) – y(3) – y(2) –

please show work, thank you!
Answer #1

IUNIE PgDn End Ins Del 5-1) gtgtyty = lt eltek + 27 + 6 2x Auxillay equation is m3 m² +m+1=0 ma (m+1)+(m+1)=0 (mati) (m+1)=0=> m= ti tre Je=c, cost to sinnt Ge tq ex - (Hx+x+x3) Jp = 040-89-0-2 14x7x + 3 = 21- (oto3_8_D) m = I1 - 9 - Je = c, cosx +> D = 1/2 c=1+30 → C= *(1-3) =* -26 -5 -6*=) → -2B = 1+5-3 > -2B = 1/2 * 2B = 3 i ri i Alw elú det elav u I I → -2A = we + -